# Quantum-limited superresolution of two incoherent point sources with unknown photon numbers

### v1

@@ Appendix C Precision for superresolution of two identical optical sources
For two incoherent optical sources with arbitrary photon numbers $N_{1}$ and $N_{2}$, the density operator for the arrival photons on the image plane can be written as
@@ C.1 With known arbitrary photon numbers
Taking the inverse of the quantum Fisher information matrix (94), the estimation precision of the centroid separation $d$ can be worked out as
@@
When the centroid separation goes to zero, the estimation separation precision $\mathcal{H}_{d}$ becomes
@@
For two Gaussian point-spread functions, the estimation precision for the centroid separation $d$ is
@@
In the limit $d\rightarrow 0$, the estimation precision of centroid separation (89) can be simplified to
@@
This recovers the results for the original quantum superresolution scheme in the seminal paper by Tsang et al. [20].
@@ C.2 Unknown arbitrary photon numbers
where $\kappa$, $\gamma$ are defined in Eq. (88) and $\delta$ is defined in Eq. (85).Taking the inverse of the quantum Fisher information matrix $\mathcal{Q}$ (94), the estimation precision of the centroid separation $d$ can be worked out as
@@
If the two photon numbers $N_{1}$, $N_{2}$ are different and $N_{2}-N_{1}$ keeps finite, the precision of centroid separation (95) can be simplified to
@@
which is of order $O(d^{2})$, where ${\rm Var}(\hat{P}^{2})=\langle\hat{P}^{4}\rangle-\langle\hat{P}^{2}\rangle^{2}$. This indicates that estimation precision for the centroid separation vanishes when $d$ is small, in drastic constraint to the above results for resolving two identical incoherent optical sources with known photon numbers. So the quantum superresolution scheme does not work and the Rayleigh’s curse takes effect again in this case, which is called the resurgence of Rayleigh’s curse.
In order to study how to avoid the resurgence of Rayleigh’s curse, we need to take a deeper insight into how the resurgence of Rayleigh’s curse occurs by expanding the numerator and denominator of $\mathcal{H}_{d}$ (95) to the second order of $d$,
@@
$\mathcal{H}_{d}=\frac{3N_{{\rm tot}}d^{2}\left(1-\epsilon^{2}\right)\langle\hat{P}^{2}\rangle{\rm Var}(\hat{P}^{2})}{12\epsilon^{2}\langle\hat{P}^{2}\rangle+d^{2}\left[3{\rm Var}(\hat{P}^{2})-4\epsilon^{2}\langle\hat{P}^{4}\rangle\right]},$ (99)
Obviously, when the photon numbers of the two incoherent point sources are the same, i.e., $\epsilon=0$, the leading constant term of the denominator of $\mathcal{H}_{d}$ vanishes and both the numerator and the denominator of $\mathcal{H}_{d}$ are of order $O(d^{2})$, so $\mathcal{H}_{d}$ keeps nonzero and independent of $d$ when $d\rightarrow 0$, which is the quantum superresolution of two incoherent point sources. In contrast, if the photon numbers of the two optical sources are different, i.e., $N_{1}\neq N_{2}$, the magnitude of the numerator of $\mathcal{H}_{d}$ is $O(d^{2})$ but the magnitude of the denominator of $\mathcal{H}_{d}$ is $O(1)$ if the difference between the two photon numbers keeps finite. In this case when $d\rightarrow 0$ the separation precision vanishes, which is the resurgence of the Rayleigh’s curse.
@@
If we further denote the ratio between the centroid separation and the total width of the two point sources as
$r=\frac{d}{2\sigma},$ (101)
@@
$\mathcal{H}_{d}=\frac{\left(1-\epsilon^{2}\right)\left({\rm e}^{r^{2}}-1-r^{2}\right)}{{\rm e}^{r^{2}}-1-r^{2}\left(1-\epsilon^{2}\right)}\mathcal{H}_{d}^{({\rm opt})},$ (102)
@@
By expanding both the numerator and denominator of Eq. (100) to the second order of $d$, one can have
@@
showing $\mathcal{H}_{d}$ keeps nonzero when $d$ is small if $\epsilon=0$ which is the quantum superresolution result and vanishes if $\epsilon$ is nonzero and keeps finite which is the resurgence of Rayleigh’s curse.
The above analysis givens an important implication that the magnitude of $\epsilon$ plays a critical role in determining the effect of the quantum superresolution. If $N_{1}$ and $N_{2}$ are different but the magnitude of $\epsilon$ is of order $O(d\sqrt{{\rm Var}(\hat{P}^{2})/\langle\hat{P}^{2}\rangle})$ or higher, the order of the denominator of $\mathcal{H}_{d}$ is still $O(d^{2})$ then and $\mathcal{H}_{d}$ will not vanish in this case. This immediately leads to the condition for the quantum superresolution with two identical incoherent optical sources in the main text.
@@ Appendix D Precision for resolving two arbitrary point sources
For two arbitrary optical point sources, the density operator can be written as
@@
When the photon numbers of the two optical sources are unknown, one needs to estimate the ratio $\epsilon$ of the difference between the two photon numbers to the total photon number in addition to the mean $\bar{X}$ and the separation $d$ of the two centroid positions. The quantum Fisher information matrix $\mathcal{Q}$ for parameters $\bar{X},\,d,\,\epsilon$ can be derived by Eq. (75) as
@@
By taking the inverse of the matrix $\mathcal{Q}$, the quantum Cramér-Rao’s bound for estimation precision of the centroid separation $d$ can be derived as
@@
For this case, the Rayleigh’s curse does not take effect and the quantum superresolution works. Actually in this case, only $A_{0}$ and $B_{0}$ are involved.
When the difference between the two point-spread functions and the difference between the two photon numbers are comparable to the magnitude of the centroid separation $d$, the higher-order terms in Eq. (114) will also contribute to $\mathcal{H}_{d}$ in the limit $d\rightarrow 0$ And the results are summarized in Table 1 of the main text.
For two Gaussian point-spread functions with centroid positions $X_{1},X_{2}$ and widths $\sigma_{1},\sigma_{2}$, the parameters $\delta,\gamma,\kappa_{1},\kappa_{2}$ can be obtained as
@@
$\kappa_{1}$ $=\frac{1}{4\sigma_{1}^{2}},\;\kappa_{2}=\frac{1}{4\sigma_{2}^{2}}.$
where $\sigma_{{\rm tot}}=\sigma_{1}+\sigma_{2}$. The estimation precision of the centroid separation $d$ turns to be
@@
If we denote
$\eta=\frac{\sigma_{2}-\sigma_{1}}{\sigma_{1}+\sigma_{2}},\;,r=\frac{d}{\sigma_{1}+\sigma_{2}},$ (120)
the estimation precision $\mathcal{H}_{d}$ (119) can be simplified as
$\mathcal{H}_{d}=\frac{\left(1-\epsilon^{2}\right)\left\{\left(1-\eta^{2}\right)\left[\left(1+\eta^{2}\right)^{2}+r^{2}\left(1-2\eta\epsilon+\eta^{2}\right)\right]-\left(1+\eta^{2}\right)^{3}{\rm e}^{\frac{r^{2}}{\eta^{2}+1}}\right\}}{\left(1-\eta^{2}\right)^{3}\left(1-\epsilon^{2}\right)r^{2}-\left(1+\eta^{2}\right)^{2}\left(1-2\eta\epsilon+\eta^{2}\right)\left[\eta^{2}+\left(\eta^{2}+1\right){\rm e}^{\frac{r^{2}}{\eta^{2}+1}}-1\right]}\mathcal{H}_{d}^{({\rm opt})},$ (121)
$\mathcal{H}_{d}^{({\rm opt})}=\frac{N}{2\sigma^{2}}.$ (122)
This implies that the ratio between the precision $\mathcal{H}_{d}$ and the optimal precision $\mathcal{H}_{d}^{({\rm opt})}$ is determined by the ratio $\epsilon$ between the difference and the total of the two photon numbers, the ratio $\eta$ between the difference and the total of the two Gaussian point-spread widths and the ratio $r$ between the centroid separation and the total of the two widths.
@@
which is Eq. (50) in the main text.
The results of the estimation precision $\mathcal{H}_{d}$ in different regimes of the photon numbers and the widths of the two point sources are summarized in Table 2 of the main text.
@@ E.1 Critical quantities
[…]
When the two point-spread functions are close to each other, $\langle\psi_{1}|\psi_{2}\rangle\rightarrow 1$, $\chi\rightarrow 0$, $2\xi-\langle\psi_{1}|\psi_{2}\rangle=2\xi-1\rightarrow 0$ , which leads to $A_{2}\rightarrow 0$, $B_{2}\rightarrow 0$, so one can immediately find that $\chi$ and $\frac{1}{2}-\xi$ are the two additional quantities that are determined by the difference between the two point-spread functions and can change the estimation precision of the centroid separation.
@@
In summary, we can identify $1-\langle\psi_{1}|\psi_{2}\rangle$, $\chi$ and $\frac{1}{2}-\xi$ as the critical quantities that are determined by the difference between the point-spread functions and will compete with each other and with the ratio $\epsilon$ beween the difference and the total of the photon number in determining the validity and the estimation precision of quantum superresolution.
@@ E.2 Relation between critical quantities
When considering the competition between the critical quantities obtained above in determining the validity of superresolution, it should be noted that those quantities are not independent of each other but closely related, and we need to find out the relation between those quantities which will impose constraints on the range of estimation precision for the centroid separation.
Assume that the difference between $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$ is small. To facilitate the following computation, we first construct an orthornormal basis for the subspace spanned by $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$,
@@
In the above equation, $u$ is a small parameter when the two point-spread functions are close to each other. The numerator ranges between $O(u^{\frac{1}{2}})$ and $O(u)$, and the denominator ranges between $O(1)$ and $O(u)$, so the order of $\chi$ is between $O(u^{1/2})$ and $O(u)$.
Similarly, for $\frac{1}{2}-\xi$, we note that

### v2

@@ Appendix C Precision for superresolution of two identical optical sources
For two incoherent optical sources if one photon is emitted, the quantum density operator for the single photon of the optical field on the image plane can be written as
@@ C.1 With known arbitrary photon numbers
Taking the inverse of the quantum Fisher information matrix (105), the estimation precision of the separation $d$ can be worked out as
@@
When the separation goes to zero, the estimation separation precision $\mathcal{H}_{d}$ becomes
@@
For two Gaussian point-spread functions, the estimation precision for the separation $d$ is
@@
In the limit $d\rightarrow 0$, the estimation precision of the separation (100) can be simplified to
@@
This recovers the results for the original superresolution scheme in the paper by Tsang et al. [20].
@@ C.2 Unknown arbitrary photon numbers
where $\kappa$, $\gamma$ are defined in Eq. (99) and $\delta$ is defined in Eq. (96).Taking the inverse of the quantum Fisher information matrix $\mathcal{Q}$ (105), the estimation precision of the separation $d$ can be worked out as
@@
If the two photon numbers $N_{1}$, $N_{2}$ are different and $N_{2}-N_{1}$ keeps finite, the precision of the separation (106) can be simplified to
@@
which is of order $O(d^{2})$, where ${\rm Var}(\hat{P}^{2})=\langle\hat{P}^{4}\rangle-\langle\hat{P}{}^{2}\rangle^{2}$. This indicates that estimation precision for the separation vanishes when $d$ is small, in drastic constraint to the above results for resolving two identical incoherent optical sources with known photon numbers. So the superresolution scheme does not work and the Rayleigh’s curse takes effect again in this case, which is sometimes called the resurgence of Rayleigh’s curse. In order to study how to address this issue, we need to take a deeper insight into the estimation precision $\mathcal{H}_{d}$ by expanding the numerator and denominator of $\mathcal{H}_{d}$ (106) to the fourth order of $d$,
@@
$\mathcal{H}_{d}=\frac{3N_{{\rm tot}}d^{2}\left(1-\epsilon^{2}\right)\langle\hat{P}{}^{2}\rangle{\rm Var}(\hat{P}^{2})}{12\epsilon^{2}\langle\hat{P}{}^{2}\rangle+d^{2}\left[3{\rm Var}(\hat{P}^{2})-4\epsilon^{2}\langle\hat{P}^{4}\rangle\right]}.$ (110)
Obviously, when the photon numbers of the two incoherent point sources are the same, i.e., $\epsilon=0$, the leading constant term of the denominator of $\mathcal{H}_{d}$ vanishes and both the numerator and the denominator of $\mathcal{H}_{d}$ are of order $O(d^{2})$, so $\mathcal{H}_{d}$ keeps nonzero and independent of $d$ when $d\rightarrow 0$, which is the quantum superresolution of two incoherent point sources. In contrast, if the photon numbers of the two optical sources are different, i.e., $N_{1}\neq N_{2}$, the magnitude of the numerator of $\mathcal{H}_{d}$ is $O(d^{2})$ but the magnitude of the denominator of $\mathcal{H}_{d}$ is $O(1)$ if the difference between the two photon numbers keeps finite. In this case when $d\rightarrow 0$ the separation precision vanishes, which is the limitation imposed by the Rayleigh’s curse.
@@
If we further define a dimensionless separation between the two Gaussian point sources as
$\widetilde{d}=\frac{d}{\sigma},$ (112)
@@
$\mathcal{H}_{d}=\frac{\left(4e^{\frac{\widetilde{d}^{2}}{4}}-\widetilde{d}^{2}-4\right)\left(1-\epsilon^{2}\right)}{4e^{\frac{\widetilde{d}^{2}}{4}}-\widetilde{d}^{2}\left(1-\epsilon^{2}\right)-4}\mathcal{H}_{d}^{({\rm opt})},$ (113)
@@
By expanding both the numerator and denominator of Eq. (111) to the fourth order of $d$, one can have
@@
where a common $d^{2}$ has been reduced from both the numerator and denominator. It shows $\mathcal{H}_{d}$ keeps nonzero when $d$ is small if $\epsilon=0$ which is the result of quantum superresolution and vanishes if $\epsilon$ is nonzero and keeps finite which is the result of Rayleigh’s criterion.
The above analysis givens an important implication that the magnitude of $\epsilon$ plays a critical role in determining the effect of superresolution. If $N_{1}$ and $N_{2}$ are different but the magnitude of $\epsilon$ is of order $O(d\sqrt{{\rm Var}(\hat{P}^{2})/\langle\hat{P}{}^{2}\rangle})$ or higher, the order of the denominator of $\mathcal{H}_{d}$ is still $O(d^{2})$ then and $\mathcal{H}_{d}$ will not vanish in this case. This immediately leads to the condition for the superresolution with two identical incoherent optical sources in the main text.
@@ Appendix D Precision for resolving two arbitrary point sources
D.1 General point-spread functions
For two arbitrary optical point sources, the density operator of a single photon can be written as
@@ D.1 General point-spread functions
When the photon numbers of the two optical sources are unknown, one needs to estimate the ratio $\epsilon$ of the difference between the two photon numbers to the total photon number in addition to the centroid $\bar{X}$ and the separation $d$ of the two sources. The quantum Fisher information matrix $\mathcal{Q}$ for parameters $\bar{X},\,d,\,\epsilon$ can be derived by Eq. (88) as
@@
By taking the inverse of the matrix $\mathcal{Q}$, the quantum Cramér-Rao’s bound for estimation precision of the separation $d$ can be derived as
@@
For this case, the Rayleigh’s criterion is violated and the superresolution works. Actually in this case, only $A_{0}$ and $B_{0}$ are involved.
When the difference between the two point-spread functions and the difference between the two photon numbers are comparable to the magnitude of the separation $d$, the higher-order terms in Eq. (125) will also contribute to $\mathcal{H}_{d}$ in the limit $d\rightarrow 0$ And the results are summarized in Table 1 of the main text.
D.2 Gaussian point-spread functions
For two Gaussian point-spread functions with positions $X_{1},X_{2}$ and widths $\sigma_{1},\sigma_{2}$, the parameters $\delta,\gamma,\kappa_{1},\kappa_{2}$ can be obtained as
@@ D.2 Gaussian point-spread functions
$\kappa_{1}$ $=\frac{1}{4\sigma_{1}^{2}},\;\kappa_{2}=\frac{1}{4\sigma_{2}^{2}},$
where $\sigma_{{\rm tot}}=\sigma_{1}+\sigma_{2}$. The estimation precision of the separation $d$ turns to be
@@
which is Eq. (51) in the main text. The results of the estimation precision $\mathcal{H}_{d}$ in different regimes of the photon numbers and the widths of the two point sources are summarized in Table 2 of the main text.
If we denote
$\eta=\frac{\sigma_{2}-\sigma_{1}}{\sigma_{1}+\sigma_{2}},\quad\widetilde{d}=\frac{d}{(\sigma_{1}+\sigma_{2})/2},$ (133)
where $\widetilde{d}$ is the dimensionless separation extended from the case of two identical point-spread functions (112), the estimation precision $\mathcal{H}_{d}$ (130) can be simplified as
$\mathcal{H}_{d}=\frac{\left(1-\epsilon^{2}\right)\left\{4\left(\eta^{2}+1\right)^{3}e^{\frac{\widetilde{d}^{2}}{4\eta^{2}+4}}-\left(1-\eta^{2}\right)\left[4\left(\eta^{2}+1\right)^{2}+\widetilde{d}^{2}\left(\eta^{2}-2\eta\epsilon+1\right)\right]\right\}}{4\left(\eta^{2}+1\right)^{2}\left(\eta^{2}-2\eta\epsilon+1\right)\left[\left(\eta^{2}+1\right)e^{\frac{\widetilde{d}^{2}}{4\eta^{2}+4}}-\left(1-\eta^{2}\right)\right]-\widetilde{d}^{2}\left(1-\eta^{2}\right)^{3}\left(1-\epsilon^{2}\right)}\mathcal{H}_{d}^{({\rm opt})},$ (134)
$\mathcal{H}_{d}^{({\rm opt})}=\frac{N}{2\sigma^{2}}.$ (135)
This implies that the ratio between the precision $\mathcal{H}_{d}$ and the optimal precision $\mathcal{H}_{d}^{({\rm opt})}$ is determined by the parameters $\epsilon$, $\eta$ and $\widetilde{d}$ only.
@@ E.1 Critical quantities
$\text{{\hbox{\kappa_{{\rm tot}}^{}}}}=\kappa_{1}+\kappa_{2}=\langle\hat{P}{}^{2}\rangle_{2}+\langle\hat{P}{}^{2}\rangle_{1}.$ (140)
When the two point-spread functions are close to each other, $\langle\psi_{1}|\psi_{2}\rangle\rightarrow 1$, $\chi\rightarrow 0$, $2\xi-\langle\psi_{1}|\psi_{2}\rangle=2\xi-1\rightarrow 0$ , which leads to $A_{2}\rightarrow 0$, $B_{2}\rightarrow 0$, so one can immediately find that $\chi$ and $\frac{1}{2}-\xi$ are the two additional quantities that are determined by the difference between the two point-spread functions and can change the estimation precision of the separation $d$.
@@
In summary, we can identify $1-\langle\psi_{1}|\psi_{2}\rangle$, $\chi$ and $\frac{1}{2}-\xi$ as the critical quantities that are determined by the difference between the point-spread functions and will compete with each other and with the ratio $\epsilon$ between the difference and the total of the photon number in determining the validity and the estimation precision of superresolution.
@@ E.2 Relation between critical quantities
When considering the competition between the critical quantities obtained above in determining the validity of superresolution, it should be noted that those quantities are not independent of each other but closely related, and we need to find out the relation between those quantities which will impose constraints on the range of estimation precision for the separation.
Assume that the difference between $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$ is small. To facilitate the following computation, we first construct an orthonormal basis for the subspace spanned by $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$,
@@
In the above equation, $u$ is a small parameter when the two point-spread functions are close to each other. The numerator ranges between $O(u^{\frac{1}{2}})$ and $O(u)$, and the denominator ranges between $O(1)$ and $O(u)$, so the order of $\chi$ is between $O(u^{1/2})$ and $O(u)$.Similarly, for $\frac{1}{2}-\xi$, we note that